OpenStudy (anonymous):

Please help, I have the answer but I do not understand how to get there. The problem is basically a kite a quadrilateral named ABCD. It says "The measure of angle B is 120 degrees and the measure of angle A is 45 degrees. What is the measure of D? The answer is 150 degrees, BUT HOW? That's what I want to know. I would really really appreciate if someone would help me! :D

OpenStudy (amistre64):

 its not a kite if d is spose to be anything other than equal to b

OpenStudy (amistre64):

or ive got the pic labeled wrong

OpenStudy (amistre64):

if a=c, and b and d are the 'tips' of the kite: the quad is spose to add up to 360 (its just 2 triangles) a+b+c+d = 360 90+120+d=360 9+12+d/10=36 d/10=15 d = 150 ....

OpenStudy (amistre64):



OpenStudy (anonymous):

Thank You, I realized what I did wrong. I didn't subtract the right numbers... I was trying to find a picture for you guys to see but you guys did it before me! Thank you thank you thank you ! :D

OpenStudy (anonymous):

the method*